# DiarizationLM: Speaker Diarization Post-Processing with Large Language Models

### v4

@@ 2 Motivating example
Prompt: In the speaker diarization transcript below, some words are potentially misplaced. Please correct those words and move them to the right speaker. Directly show the corrected transcript without explaining what changes were made or why you made those changes. <spk:1> Good morning Patrick, how <spk:2> are you? Good, good. How are you Tom? Pretty <spk:1> good. Going to work? <spk:2> Yes. Busy day. How are your kids? Do they go <spk:1> to school? Oh they are too young for that. I sent them to daycare earlier <spk:2> today. Oh yeah I forgot about that. Completion: <spk:1> Good morning Patrick, how are you? <spk:2> Good, good. How are you Tom? <spk:1> Pretty good. Going to work? <spk:2> Yes. Busy day. How are your kids? Do they go to school? <spk:1> Oh, they are too young for that. I sent them to daycare earlier today. <spk:2> Oh yeah, I forgot about that.
@@ 3.2 Prompt builder
Word sequence: ["good", "morning", "how", "are", "you"] Speaker sequence: [1, 1, 2, 2, 2] Text representation: "<spk:1> good morning <spk:2> how are you"
@@ 4.2 Metrics
All three metrics reported in this paper (WER, WDER, and cpWER) are macro metrics, i.e. both numerators and denominators are aggregated on the entire dataset.
@@ 4.3 Models
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix “ --> ” (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix “ [eod]” (short for “end of document”; note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
@@ A.1 Autofilling speaker names
Prompt: [video title] Tom meets Patrick in the parking lot [video description] A short conversation Replace the <spk:?> tokens from the below video caption with actual person names. <spk:1> Good morning Patrick, how are you? <spk:2> Good, good. How are you Tom? <spk:1> Pretty good. Going to work? <spk:2> Yes. Busy day. How are your kids? Do they go to school? <spk:1> Oh they are too young for that. I sent them to daycare earlier today. <spk:2> Oh yeah I forgot about that. Completion: Based on the video title "Tom meets Patrick in the parking lot", here’s the transcript with speaker names replaced: Tom: Good morning Patrick, how are you? Patrick: Good, good. How are you Tom? Tom: Pretty good. Going to work? Patrick: Yes. Busy day. How are your kids? Do they go to school? Tom: Oh they are too young for that. I sent them to daycare earlier today. Patrick: Oh yeah, I forgot about that.
@@ A.2 Autofilling speaker roles
Prompt: Replace the <spk:?> tokens from the below conversation with roles like teacher, student, doctor, patient, etc. <spk:1> Hi, how can I help you today? <spk:2> Hi, I recently often feel quite dizzy at work. <spk:1> Do you have any of these symptoms? Like coughing, fever, or running nose? <spk:2> No, I don’t think so. <spk:1> What do you usually eat for breakfast, lunch and dinner? <spk:2> I’m on a diet, so not eating very much. Mostly just one slice of toast in the morning, and eating salad for lunch and dinner. Completion: Doctor: Hi, how can I help you today? Patient: Hi, I recently often feel quite dizzy at work. Doctor: Do you have any of these symptoms? Like coughing, fever, or running nose? Patient: No, I don’t think so. Doctor: What do you usually eat for breakfast, lunch, and dinner? Patient: I’m on a diet, so not eating very much. Mostly just one slice of toast in the morning, and eating salad for lunch and dinner.
@@ A.3 Replacing the orchestration module
Prompt: Here we define the problem of speaker-transcript alignment. The transcript is represented by multiple entries of text, where each text has a starting time and an ending time. The speaker is also represented in this format. The alignment problem will assign a speaker to each word in the text, based on which speaker overlaps the most with that word in time. Below is an example. Transcript represented in format "[start - end] text": [0 - 2.3] Good morning Patrick [2.5 - 5.2] how are you? [5.6 - 6.1] Good, good. [6.2 - 8.3] How are you Tom? [9.2 - 9.9] Pretty good. [10.0 - 11.1] Going to work? [12.5 - 13.6] Yes. Busy day. Speaker represented in format "[start - end] speaker": [0 - 5.1] <spk:1> [5.3 - 8.7] <spk:2> [9.2 - 10.9] <spk:1> [12.1 - 13.5] <spk:2> The alignment result should be: [0 - 2.3] <spk:1> Good morning Patrick [2.5 - 5.2] <spk:1> how are you? [5.6 - 6.1] <spk:2> Good, good. [6.2 - 8.3] <spk:2> How are you Tom? [9.2 - 9.9] <spk:1> Pretty good. [10.0 - 11.1] <spk:1> Going to work? [12.5 - 13.6] <spk:2> Yes. Busy day. Based on the example above, we solve the alignment problem below. Transcript represented in format "[start - end] text": [0 - 5.6] Hi, how can I help you today? [6.2 - 11.1] Hi, I recently often feel quite dizzy at work. [11.6 - 15.5] Do you have any of these symptoms? [16.6 - 18.5] Like coughing, fever, or running nose? [20.0 - 21.1] No, I don’t think so. [22.2 - 29.9] What do you usually eat for breakfast, lunch and dinner? [31.2 - 34.8] I’m on a diet, so not eating very much. [35.2 - 39.8] Mostly just one slice of toast in the morning, and eating salad for lunch and dinner. Speaker represented in format "[start - end] speaker": [0.3 - 5.3] <spk:1> [6.0 - 12.0] <spk:2> [12.9 - 20.1] <spk:1> [20.2 - 21.0] <spk:2> [21.8 - 31.1] <spk:1> [32.4 - 40.7] <spk:2> The alignment result should be: Completion: Here’s the aligned transcript: [0 - 5.6] <spk:1> Hi, how can I help you today? [6.2 - 11.1] <spk:2> Hi, I recently often feel quite dizzy at work. [11.6 - 15.5] <spk:1> Do you have any of these symptoms? [16.6 - 18.5] <spk:1> Like coughing, fever, or running nose? [20.0 - 21.1] <spk:2> No, I don’t think so. [22.2 - 29.9] <spk:1> What do you usually eat for breakfast, lunch and dinner? [31.2 - 34.8] <spk:2> I’m on a diet, so not eating very much. [35.2 - 39.8] <spk:2> Mostly just one slice of toast in the morning, and eating salad for lunch and dinner. To align the transcript and speaker segments, we assign each word in the transcript to the speaker whose segment overlaps the most with that word in time.

### v5

@@ 2 Motivating example
Prompt:
In the speaker diarization transcript below, some words are potentially misplaced. Please correct those words and move them to the right speaker. Directly show the corrected transcript without explaining what changes were made or why you made those changes.
<spk:1> Good morning Patrick, how
<spk:2> are you? Good, good. How are you Tom? Pretty
<spk:1> good. Going to work?
<spk:2> Yes. Busy day. How are your kids? Do they go
<spk:1> to school? Oh they are too young for that. I sent them to daycare earlier
<spk:2> today. Oh yeah I forgot about that.
Completion:
<spk:1> Good morning Patrick, how are you?
<spk:2> Good, good. How are you Tom?
<spk:1> Pretty good. Going to work?
<spk:2> Yes. Busy day. How are your kids? Do they go to school?
<spk:1> Oh, they are too young for that. I sent them to daycare earlier today.
<spk:2> Oh yeah, I forgot about that.
@@ 3.2 Prompt builder
Word sequence: ["good", "morning", "how", "are", "you"]
Speaker sequence: [1, 1, 2, 2, 2]
Text representation: "<spk:1> good morning <spk:2> how are you"
@@ 4.2 Metrics
All three metrics reported in this paper (WER, WDER, and cpWER) are micro metrics, i.e. both numerators and denominators are aggregated on the entire dataset.
@@ 4.3 Models
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix “␣-->␣” (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix “␣[eod]” (short for “end of document”; note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
@@ A.1 Autofilling speaker names
Prompt:
[video title] Tom meets Patrick in the parking lot
[video description] A short conversation
Replace the <spk:?> tokens from the below video caption with actual person names.
<spk:1> Good morning Patrick, how are you?
<spk:2> Good, good. How are you Tom?
<spk:1> Pretty good. Going to work?
<spk:2> Yes. Busy day. How are your kids? Do they go to school?
<spk:1> Oh they are too young for that. I sent them to daycare earlier today.
<spk:2> Oh yeah I forgot about that.
Completion:
Based on the video title "Tom meets Patrick in the parking lot", here’s the transcript with speaker names replaced:
Tom: Good morning Patrick, how are you?
Patrick: Good, good. How are you Tom?
Tom: Pretty good. Going to work?
Patrick: Yes. Busy day. How are your kids? Do they go to school?
Tom: Oh they are too young for that. I sent them to daycare earlier today.
Patrick: Oh yeah, I forgot about that.
@@ A.2 Autofilling speaker roles
Prompt:
Replace the <spk:?> tokens from the below conversation with roles like teacher, student, doctor, patient, etc.
<spk:1> Hi, how can I help you today?
<spk:2> Hi, I recently often feel quite dizzy at work.
<spk:1> Do you have any of these symptoms? Like coughing, fever, or running nose?
<spk:2> No, I don’t think so.
<spk:1> What do you usually eat for breakfast, lunch and dinner?
<spk:2> I’m on a diet, so not eating very much. Mostly just one slice of toast in the morning, and eating salad for lunch and dinner.
Completion:
Doctor: Hi, how can I help you today?
Patient: Hi, I recently often feel quite dizzy at work.
Doctor: Do you have any of these symptoms? Like coughing, fever, or running nose?
Patient: No, I don’t think so.
Doctor: What do you usually eat for breakfast, lunch, and dinner?
Patient: I’m on a diet, so not eating very much. Mostly just one slice of toast in the morning, and eating salad for lunch and dinner.
@@ A.3 Replacing the orchestration module
Prompt:
Here we define the problem of speaker-transcript alignment. The transcript is represented by multiple entries of text, where each text has a starting time and an ending time. The speaker is also represented in this format. The alignment problem will assign a speaker to each word in the text, based on which speaker overlaps the most with that word in time.
Below is an example.
Transcript represented in format "[start - end] text":
[0 - 2.3] Good morning Patrick
[2.5 - 5.2] how are you?
[5.6 - 6.1] Good, good.
[6.2 - 8.3] How are you Tom?
[9.2 - 9.9] Pretty good.
[10.0 - 11.1] Going to work?
[12.5 - 13.6] Yes. Busy day.
Speaker represented in format "[start - end] speaker":
[0 - 5.1] <spk:1>
[5.3 - 8.7] <spk:2>
[9.2 - 10.9] <spk:1>
[12.1 - 13.5] <spk:2>
The alignment result should be:
[0 - 2.3] <spk:1> Good morning Patrick
[2.5 - 5.2] <spk:1> how are you?
[5.6 - 6.1] <spk:2> Good, good.
[6.2 - 8.3] <spk:2> How are you Tom?
[9.2 - 9.9] <spk:1> Pretty good.
[10.0 - 11.1] <spk:1> Going to work?
[12.5 - 13.6] <spk:2> Yes. Busy day.
Based on the example above, we solve the alignment problem below.
Transcript represented in format "[start - end] text":
[0 - 5.6] Hi, how can I help you today?
[6.2 - 11.1] Hi, I recently often feel quite dizzy at work.
[11.6 - 15.5] Do you have any of these symptoms?
[16.6 - 18.5] Like coughing, fever, or running nose?
[20.0 - 21.1] No, I don’t think so.
[22.2 - 29.9] What do you usually eat for breakfast, lunch and dinner?
[31.2 - 34.8] I’m on a diet, so not eating very much.
[35.2 - 39.8] Mostly just one slice of toast in the morning, and eating salad for lunch and dinner.
Speaker represented in format "[start - end] speaker":
[0.3 - 5.3] <spk:1>
[6.0 - 12.0] <spk:2>
[12.9 - 20.1] <spk:1>
[20.2 - 21.0] <spk:2>
[21.8 - 31.1] <spk:1>
[32.4 - 40.7] <spk:2>
The alignment result should be:
Completion:
Here’s the aligned transcript:
[0 - 5.6] <spk:1> Hi, how can I help you today?
[6.2 - 11.1] <spk:2> Hi, I recently often feel quite dizzy at work.
[11.6 - 15.5] <spk:1> Do you have any of these symptoms?
[16.6 - 18.5] <spk:1> Like coughing, fever, or running nose?
[20.0 - 21.1] <spk:2> No, I don’t think so.
[22.2 - 29.9] <spk:1> What do you usually eat for breakfast, lunch and dinner?
[31.2 - 34.8] <spk:2> I’m on a diet, so not eating very much.
[35.2 - 39.8] <spk:2> Mostly just one slice of toast in the morning, and eating salad for lunch and dinner.
To align the transcript and speaker segments, we assign each word in the transcript to the speaker whose segment overlaps the most with that word in time.